OpenStudy (anonymous):

Which correctly describes the concavity of the function f(x) = x^3 – 5x + 7?

OpenStudy (amistre64):

which which?

OpenStudy (bahrom7893):

3x^2 - 5 = 0 x^2 = 5/3 x = + or - 5/3

OpenStudy (bahrom7893):

now: 6x, is f '' so when is that positive?

OpenStudy (bahrom7893):

-infinity to -1.666, -1.666 to +1.666 or 1.666 to infinity?

OpenStudy (anonymous):

is the concave downward or upward though?

OpenStudy (amistre64):

a cubic has 2 concavities; one is cave down the other is cave up; just depends on where you are at on it

OpenStudy (amistre64):

your question inplies that there are options to choose from

OpenStudy (bahrom7893):

so: try -2: 6*(-2) = -12, therefore conc down on (-inf; 1.66)

OpenStudy (bahrom7893):

i meant (-inf; -1.6)

OpenStudy (bahrom7893):

6*1 = 6, therefore concave up on (-1.6,1.6)

OpenStudy (anonymous):

what about (-3.0)? that would be 6 so concave upward? Am I on the right track?

OpenStudy (bahrom7893):

No where did u get (-3,0) from?

OpenStudy (anonymous):

one of the question answers. lol. along with downward interval at 0,3 upward at 0,3 or downward at-3,3.Those are just some question choices. I wanted to learn how to do it so i could figure out how to plug in the answers myself. lmao

OpenStudy (bahrom7893):

okay look just take a number from any of those intervals, like say x=-2 which is inside -3,0 and plug that into 6x. If it's negative then it's concave down, positive - conc up

OpenStudy (anonymous):

so from the point -3,3 the answer is -18,18 so the concave is downward?

OpenStudy (amistre64):

sorry, i still cant make out what it is your asking for :/

OpenStudy (anonymous):

arggg. This is the question: Which correctly describes the concavity of the function f(x) = x^3 – 5x + 7? concave upward in the interval (-3, 0) concave downward in the interval (0, 3) concave upward in the interval (0, 3) concave downward in the interval (-3, 3)

OpenStudy (anonymous):

Those are the different options. I didnt know those answers were necessary. I wanted to try to figure it out on my own. Lol